### v1

# ContextGen: Contextual Layout Anchoring for Identity-Consistent Multi-Instance Generation

Abstract
Multi-instance image generation (MIG) remains a significant challenge for modern diffusion models due to key limitations in achieving precise control over object layout and preserving the identity of multiple distinct subjects. To address these limitations, we introduce ContextGen, a novel Diffusion Transformer framework for multi-instance generation that is guided by both layout and reference images. Our approach integrates two key technical contributions: a Contextual Layout Anchoring (CLA) mechanism that incorporates the composite layout image into the generation context to robustly anchor the objects in their desired positions, and Identity Consistency Attention (ICA), an innovative attention mechanism that leverages contextual reference images to ensure the identity consistency of multiple instances. Recognizing the lack of large-scale, hierarchically-structured datasets for this task, we introduce IMIG-100K, the first dataset with detailed layout and identity annotations. Extensive experiments demonstrate that ContextGen sets a new state-of-the-art, outperforming existing methods in control precision, identity fidelity, and overall visual quality. Our project page is at https://nenhang.github.io/ContextGen/.
Figure 1: Representative showcases of our work. Upper panel: Our multi-subject-driven generation results versus existing open-source SOTA (OmniGen2) and proprietary models (Nano Banana, GPT-4o). Lower panel: Our layout-to-image generation examples using different layouts.
1 Introduction
Diffusion-based models (Ho et al., 2020) have significantly expanded the horizons of image customization, with many recent systems (e.g., FLUX (Labs, 2024b)) adopting the Diffusion Transformer (DiT) (Peebles & Xie, 2022) framework for its enhanced generation quality. Recent developments in subject-driven image generation, such as OmniGen2 (Wu et al., 2025b), and layout-to-image synthesis, exemplified by MS-Diffusion (Wang et al., 2025), have further broadened the scope of customization, enabling control over both content and composition in generated images.
However, current methods face three fundamental limitations: (1) Inadequate position control, where existing layout guidance fails to achieve accurate spatial precision for user-specified arrangements; (2) Weak identity preservation, as subject-driven approaches struggle to maintain fine details across multiple instances, particularly with an increasing number of reference images. (3) Lack of high-quality training data, as existing datasets do not provide large-scale, precisely aligned pairs of reference images and layout annotations for multi-instance scenarios. These deficiencies collectively hinder the simultaneous achievement of compositional accuracy and identity fidelity.
To address these challenges, we propose ContextGen, a novel DiT-based framework that enables multi-instance generation by unifying two key modalities. First, we use a composite layout image for precise spatial control. As shown in the setup stage of Fig. 2, this layout image can be either user-provided or automatically synthesized. Second, we integrate reference images to overcome the limitations of layout-only generation, such as instance information loss due to overlaps and dimensional compression. By incorporating these modalities into a unified contextual framework, ContextGen achieves both precise spatial control and high instance-level identity consistency.
Our work introduces three key innovations and contributions: (1) Contextual Layout Anchoring (CLA), which leverages contextual learning to anchor each instance at its desired position by incorporating the layout image into the generation context, thereby achieving robust layout control; and (2) Identity Consistency Attention (ICA), a novel attention mechanism which propagates fine-grained information from contextual reference images to their respective desired locations, thereby preserving the identity of multiple instances. Complementing these mechanisms is an enhanced position indexing strategy that systematically organizes and differentiates multi-image relationships. (3) A large-scale, hierarchically-structured dataset, IMIG-100K, which we curate with annotated bounding boxes and identity-matched references to directly address the current data scarcity in Image-guided Multi-instance Image Generation, with hierarchical samples shown in Fig. 3.
Our method achieves state-of-the-art performance across three benchmarks. On (1) COCO-MIG (Zhou et al., 2024b), it improves instance-level success rate by +3.4% and spatial accuracy (mIoU) by +5.9% over prior art. For (2) LayoutSAM-Eval (Zhang et al., 2024), it attains the highest scores in texture and color fidelity, demonstrating superior detail preservation. Most notably, on (3) LAMICBench++ (Chen et al., 2025b), our approach outperforms all open-source models by +1.3% average score and even surpasses commercial systems like GPT-4o in identity retention (+13.3%). These gains validate CLA’s layout robustness and ICA’s effectiveness in multi-instance scenarios.
In summary, our key contributions are as follows:
  • ContextGen: A novel DiT-based framework with Contextual Layout Anchoring (CLA) for robust layout control and Identity Consistency Attention (ICA) for precise identity preservation.
  • IMIG-100K: The first large-scale, hierarchically-structured image-guided multi-instance-generation dataset with layout and identity annotations.
  • SOTA Performance: We achieve state-of-the-art results, outperforming existing methods in layout control, identity preservation, and visual quality.
Figure 2: Overview of ContextGen. Left (Setup Stage): Options to composite the Layout Image. Middle (Model Core): Central generation architecture using FLUX DiT-Blocks. Right (Attention Mechanisms): Details of $M$-Attention components (Position Indexing, CLA and ICA).
2 Related Work
2.1 Diffusion Models
Diffusion models have evolved from UNet architectures (Ho et al., 2020; Rombach et al., 2022) to transformer-based approaches like DiT (Peebles & Xie, 2022), enabling scalable multimodal generation as seen in Stable Diffusion 3 (Esser et al., 2024). FLUX (Labs, 2024b) further advanced this by unifying visual and textual inputs through multi-modal attention mechanism.
2.2 Instance-level Controllable Image Generation
GLIGEN (Li et al., 2023) pioneered the layout-to-image generation paradigm. Follow-up studies utilizing UNet-based methods like InstanceDiffusion (Wang et al., 2024) and MIGC (Zhou et al., 2024a), or DiT-based approaches like EliGen (Zhang et al., 2025a) and 3DIS  (Zhou et al., 2024c), have demonstrated enhanced capabilities in handling multiple instances. Current state-of-the-art frameworks like OmniGen2 (Wu et al., 2025b) and DreamO (Mou et al., 2025) process multi-subject conditions via integrated token sequences but face identity degradation with many subjects. While MS-Diffusion (Wang et al., 2025) and LAMIC (Chen et al., 2025b) combine reference-driven generation with layout control, challenges remain in layout precision and identity consistency.
3 Method
3.1 Preliminaries
Multimodal Diffusion Transformers (MM-DiT)
Recent architectures have replaced modality-specific cross-attention with unified multimodal processing. The MM-Attention operation concatenates image tokens $\mathbf{t}_{\text{image}}$ and text embeddings $\mathbf{t}_{\text{text}}$ into a single sequence $\mathbf{T}=[\mathbf{t}_{\text{text}},\mathbf{t}_{\text{image}}]$, enabling joint self-attention across modalities. Stable Diffusion 3/3.5 (Esser et al., 2024; stability.ai, 2024) and FLUX (Labs, 2024b), treats all modalities within a shared latent space. The framework naturally supports in-context learning by allowing arbitrary interleaving of visual and textual tokens, while maintaining stable gradient flow across modalities during end-to-end training.
Position Indexing and Attention Mask in MM-Attention
To address the permutation-invariance of the Transformer architecture, Rotatory Position Embedding (RoPE) (Su et al., 2023) was introduced to encode relative positional information. Adapting this for a unified multimodal space, the FLUX.1-Dev architecture proposes a novel extension of RoPE that employs a ternary position encoding scheme. This scheme assigns a position index $\mathbf{p}_{i}=(m,i,j)$ to each token in the sequence. The first component $m$ is set to 0 and is retained for further use. For text tokens, the spatial coordinates $(i,j)$ are fixed at $(0,0)$, while for image tokens, they correspond to the spatial coordinates $(i,j)$ in the 2D noise latent space. This set of position indices $\{\mathbf{p}_{i}\}$ for the sequence forms a position index matrix $\mathbf{P}$.
The unified attention mechanism, which is controlled by the attention mask $\mathbf{M}$, integrates this positional information through RoPE. Specifically, the rotation matrix $\mathbf{R}$ is computed by applying the RoPE formulation to the position index matrix $\mathbf{P}$, a process we denote as $\mathbf{R}=\text{Rotate}(\mathbf{P})$. This resulting matrix $\mathbf{R}$ is then utilized to apply a rotation to the query ($\mathbf{Q}$) and key ($\mathbf{K}$) embeddings before the dot-product calculation. The attention is then calculated as:
$\text{MM-Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{(\mathbf{R}\mathbf{Q})(\mathbf{R}\mathbf{K})^{\top}}{\sqrt{d}}\odot\mathbf{M}\right)\mathbf{V}$ (1)
The symbol $\odot$ denotes element-wise multiplication. In the FLUX.1 series, a self-attention mechanism is employed where queries, keys, and values are all derived from the unified token sequence $\mathbf{T}$, and $\mathbf{M}$ is an all-True matrix, enabling full attention across all tokens.
3.2 Contextual Attention with Layout Anchoring and Identity Preservation
Contextual Conditioning with Layout and Reference Images
Recent studies in image-to-image (I2I) tasks have demonstrated the effectiveness of using a diptych, a side-by-side reference image pair, to guide diffusion models (Shin et al., 2024; Song et al., 2025; Zhang et al., 2025b). Building upon this, our framework introduces a novel layout control strategy by integrating a composite layout image into the generation context. This can be done either by a user-defined composition, which offers greater control and is often more aligned with specific user intent, or by our automated sorting algorithm (mentioned in Sec. A.2) based on the occlusion ratio of all instances. This composite diptych serves as the primary input for our Contextual Layout Anchoring (CLA) mechanism, which is designed to enforce a robust spatial structure by anchoring objects to their desired locations.
However, relying solely on this composite layout image presents two major challenges. First, the process of compositing multiple high-resolution instances into a single image results in a compressed representation that leads to a loss of fine-grained details. Second, in scenarios with significant instance overlap, the process of compositing may result in information loss or detail degradation. To address these issues, we integrate the original, high-fidelity reference images alongside the diptych, inspired by subject-driven generation techniques (Wu et al., 2025c; Mou et al., 2025). Then the unified token sequence $\mathbf{T}$ mentioned in Sec. 3.1 is constructed as:
$\mathbf{T}=[\mathbf{t}_{\text{text}},\mathbf{t}_{\text{image}},\mathbf{t}_{\text{layout}},\mathbf{t}_{\text{ref}_{1}},\cdots,\mathbf{t}_{\text{ref}_{N}}]$ (2)
Our Identity Consistency Attention (ICA) mechanism incorporates these tokenized reference images $\{\mathbf{t}_{\text{ref}_{i}}\}$ into the context to preserve instance-specific attributes and details, effectively mitigating the issues of detail loss in overlapping regions, thereby ensuring a complementary relationship between the robust layout guidance from CLA and the precise detail preservation from ICA.
Contextual Layout Anchoring (CLA)
Inspired by the functional specialization observed in DiT layers (Zhou et al., 2025; Zhang et al., 2024), we propose a hierarchical attention architecture to process the unified token sequence. As shown in the middle panel of Fig. 2, the CLA mechanism operates in the front and back layers, focusing primarily on global context and structural composition. The CLA mask, detailed in the right panel of Fig. 2, ensures broad communication across the text, image, and layout modalities. Using the token sets defined ($\mathcal{T}\!=\!\{\mathbf{t}_{\text{text}}\}$, $\mathcal{I}\!=\!\{\mathbf{t}_{\text{image}}\}$, $\mathcal{L}\!=\!\{\mathbf{t}_{\text{layout}}\}$, and $\mathcal{R}_{n}\!=\!\{\mathbf{t}_{\text{ref}_{n}}\}$) and reference bounding boxes $\{B_{n}\}_{n=1}^{N}$, the attention mask for CLA is defined as:
$\text{M}_{\text{CLA}}(q,k)=\begin{cases}1&\text{if }q\in\mathcal{T}\cup\mathcal{I}\cup\mathcal{L}\text{ and }k\in\mathcal{T}\cup\mathcal{I}\cup\mathcal{L}\\ &\text{or }q\in\mathcal{R}_{n}\text{ and }k\in\mathcal{T}\cup\mathcal{R}_{n}\\ 0&\text{otherwise}\end{cases}$ (3)
where $q$ and $k$ are arbitrary tokens from the query and key sequences, respectively.
Identity Consistency Attention (ICA)
While the front and back layers perform global spatial anchoring, we introduce the ICA mechanism in the middle layers to facilitate detailed, instance-level identity injection. As detailed in the right panel of Fig. 2, ICA operates by applying a specialized attention mask, $\text{M}_{\text{ICA}}$, for tokens located within a specific bounding box. For a query token $q\in B_{n}$, the attention mask is defined as:
$\text{M}_{\text{ICA}}(q,k)=\begin{cases}1&\text{if }k\in\mathcal{T}\cup B_{n}\cup\mathcal{R}_{n}\\ 0&\text{otherwise}\\ \end{cases}\qquad\text{if }q\in B_{n}$ (4)
The core function of $\text{M}_{\text{ICA}}$ is the forced connection between $q$ and its corresponding reference tokens $\mathcal{R}_{n}$, ensuring reliable identity transfer. Tokens outside any bounding box (i.e., background) default to the mask used by CLA. This hierarchical strategy effectively transitions our framework from global layout control to refined instance-level identity preservation.
Instance-Wise Position Indexing
The ternary position encoding scheme described in Sec. 3.1 was extended in FLUX.1-Kontext (Labs et al., 2025) to handle image editing, where the first component of the position index, $m$, was set to 1 for edit tokens. Inspired by this work and other existing work (Wu et al., 2025c) that shows providing distinct and non-overlapping position indices for each image sequence significantly improves the model’s ability to differentiate between various images, we propose a refined position encoding strategy to systematically structure the relationships within our unified token sequence $\mathbf{T}$ (Eq. 2).
  • Basic Part: The primary noise latent $\mathbf{t}_{\text{image}}$ retains the original $(0,i,j)$ indexing, ensuring spatial coherence within the target image.
  • Auxiliary Part: Tokens from auxiliary inputs, including layout image and reference images, are assigned a unique index. They are indexed as $(1,W_{n}+i,H_{n}+j)$, where $W_{n}=\sum_{k=1}^{n-1}w_{k}$ and $H_{n}=\sum_{k=1}^{n-1}h_{k}$ are cumulative offsets aggregating the dimensions of all preceding conditioning images. This guarantees unique positional identifiers for each conditioning image, even when they are concatenated.
This approach allows the attention mechanism to distinguish between tokens from the noise latent and auxiliary inputs, as well as to differentiate between tokens from various conditioning images.
3.3 IMIG-100K: An Image-Guided Multi-Instance-Generation Dataset
High-fidelity image-guided multi-instance generation is severely limited by the lack of suitable training data. While existing large-scale datasets  (Lin et al., 2015; Deng et al., 2009) provide diverse instances, they often lack the aesthetic quality and annotation granularity required for modern diffusion models. Conversely, recent subject-driven datasets (Tan et al., 2025; Xiao et al., 2024) exhibit high visual quality but are limited by their low instance multiplicity per image. To bridge this gap, we introduce IMIG-100K, a new large-scale dataset created using the FLUX framework (Labs, 2024b). This dataset is specifically designed to support multi-instance generation by providing high-resolution, high-fidelity data with precise layout and reference images.
Dataset Structure and Key Features
To robustly train the diverse capabilities required for identity-consistent multi-instance generation, the IMIG-100K dataset is systematically structured into three specialized sub-datasets. These subsets collectively facilitate the comprehensive training of our framework, with examples shown in Fig. 3.
Figure 3: Image Samples of IMIG-100K Dataset.
  1. 1.
    Basic Instance Composition (50K samples): This subset focuses on foundational compositional skills. The ground truth images are generated by the text-to-image model FLUX.1-Dev  (Labs, 2024b), and we derive reference images using detection and segmentation models (Liu et al., 2023; Ravi et al., 2024; Dai et al., 2025). These reference images undergo minimal post-processing, including basic lighting adjustments.
  2. 2.
    Complex Instance Interaction (50K samples): Designed for more complex scenarios with up to 8 instances per image, this subset’s data construction is similar to the basic part. However, the reference images are semantically edited to simulate real-world interactions, including occlusion, viewpoint rotation, and object pose changes.
  3. 3.
    Flexible Composition with References (10K samples): Unlike the previous two subsets, this unique subset is designed to train the model’s robustness in handling low-consistency inputs. We first generate individual reference instances using the FLUX.1-Dev model. These are then composited into ground truth scenes by subject-driven models (Wu et al., 2025c; Mou et al., 2025), allowing for a much greater degree of flexibility and transformation in the composited instances relative to their original references. A key step involves rigorous filtering to ensure identity consistency from the references (Guo et al., 2021; Oquab et al., 2023).
All textual prompts are generated by advanced large language models (DeepSeek-AI, 2025; Comanici et al., 2025; OpenAI, 2024), ensuring diverse and high-quality descriptions.
4 Experiments
4.1 Experimental Setting
Training Details
We initialize the model with FLUX.1-Kontext (Labs et al., 2025) without introducing additional parameters and fine-tune it using LoRA (Low-Rank Adaptation) (Hu et al., 2021) with LoRA Rank 512. We perform training on 4× NVIDIA A100 GPUs with a total batch size of 16. The model is tuned on the three hierarchical sub-datasets described in Sec. 3.3 for 5K steps, employing the Prodigy optimizer (Mishchenko & Defazio, 2024) with its default learning rate. We also employ Direct Preference Optimization (DPO) (Rafailov et al., 2024) to refine text-visual alignment and user preference. These enable the model to evolve from mastering simple compositions to synthesizing complex multi-instance scenes.
Benchmark Datasets
We employ three distinct benchmark datasets for evaluation.
(1) LAMICBench++: A specialized benchmark for evaluating identity preservation and feature consistency in subject-driven generation. We extend the multi-image composition benchmark from LAMICBench (Chen et al., 2025b), aggregating multi-category reference images (humans, animals, objects, etc.) from established datasets including XVerseBench (Chen et al., 2025a), DreamBench++(Peng et al., 2025) and MS-Bench (Wang et al., 2025). In particular, we construct a dataset of 160 cases in total, including 50 cases with 2 reference images, 40 with 3, 30 with 4, 20 with 5, and 20 with over 5 reference images. These cases are divided into two categories: Fewer Subjects ($\leq 3$ reference images) and More Subjects ($\geq 4$ reference images). In this benchmark, we adapt and slightly modified the four evaluation metrics from the original work: (1) Global text-image consistency (ITC) evaluated through visual-question-answering (VQA)  (Ye et al., 2024), with approximately 2K questions (4-12 per item); (2) Object preservation (IPS) (Liu et al., 2023; Oquab et al., 2023); (3) Facial identity retention (IDS) (Guo et al., 2021); (4) Aesthetic quality (AES) (Schuhmann, 2023).
(2) COCO-MIG (Zhou et al., 2024b): A benchmark designed to evaluate spatial and attribute accuracy in layout-to-image generation, comprising 800 images from COCO Dataset (Lin et al., 2015) with color-annotated instances. The evaluation metrics include: (1) Global and instance level success rate (SR and I-SR) determined by spatial accuracy (mIoU) and color correctness; (2) Multi-scale semantic consistency through global and local CLIP Scores (G-C and L-C).
(3) LayoutSAM-Eval (Zhang et al., 2024): A open-set benchmark for layout-to-image evaluation, featuring 5K prompts with exhaustive entity-level annotations, from which we filter 1K samples with sufficiently large bounding boxes for reliable instance evaluation. We adapt the original work’s metrics: (1) Fine-grained entity accuracy (spatial, color, textural, shape) evaluated using MLLM (Yao et al., 2024); (2) Holistic quality metrics: CLIP Score for semantic alignment and Pick Score (Kirstain et al., 2023) for human preference.
4.2 Baselines
We compare our method against a comprehensive set of state-of-the-art baselines across relevant domains. For layout-to-image generation, we include pioneering works such as LAMIC (Chen et al., 2025b) and MS-Diffusion (Wang et al., 2025). To evaluate spatial control, we also benchmark against CreatiLayout (Zhang et al., 2024), EliGen (Zhang et al., 2025a), MIGC (Zhou et al., 2024b), InstanceDiffusion (Wang et al., 2024), and GLIGEN (Li et al., 2023). In the domain of subject-driven generation, we benchmark against OmniGen2 (Wu et al., 2025b), DreamO (Mou et al., 2025), UNO (Wu et al., 2025c), XVerse (Chen et al., 2025a), and MIP-Adapter (Huang et al., 2024) to specifically assess identity preservation. For a cutting-edge benchmark, we highlight the latest proprietary models, including Nano Banana (formally named Gemini 2.5 Flash Image, Google’s latest multimodal model) (DeepMind, 2025) and GPT-4o-Image (OpenAI, ), as well as leading open-source models like Qwen-Image-Edit (Wu et al., 2025a) and FLUX.1-Kontext (Labs et al., 2025).
Table 1: Quantitative results on LAMICBench++. Performance rankings: bold (highest), underline (second highest), wavy underline (third highest). The benchmark provides complete manual annotations for all method requirements: layout-aware methods () use our pre-annotated bounding boxes, while single-image-editing methods () use our manually composited layout images.
Figure 4: Qualitative results on LAMICBench++.
Table 2: Quantitative results on COCO-MIG and LAMICBench++. Image-guided methods () use our pre-generated images by FLUX.1-Dev (Labs, 2024b)
Figure 5: Qualitative results on COCO-MIG. We use red dashed box to indicate the missing, merged, dislocated and incorrectly attributed instances.
Figure 6: Qualitative results on LayoutSam-Eval.
4.3 Comparison
Identity Preservation and Overall Quality
Quantitative results on LAMICBench++ in Tab. 1 show that our method excels in object preservation and facial identity retention. In Fewer Subjects, we achieve the highest IPS with competitive IDS. This advantage amplifies in More Subjects, while other open-source models experience significant drops in these metrics. Compared to closed-source models (GPT-4o and Nano Banana), we show a strategic trade-off: while slightly trailing in ITC and AES, we outperform them significantly in both IPS and IDS. This balanced performance yields our superior overall benchmark score (64.66 vs 63.71/64.11), demonstrating exceptional capability in preserving both objects and identities simultaneously.
Fig. 4 demonstrates our method’s superior performance in preserving both content and style across diverse scenarios. Our approach consistently maintains accurate object relationships and fine details where other methods fail - evident in the precise rendering of facial identities (old man’s wrinkles), object features (shape of the vase, appearance of piggy bank, color and texture of Sphynx cat).
Layout Control and Attribute Binding
Tab. 2 shows our method achieves superior layout control with the highest correctness on COCO-MIG. Direct comparison with text-guided L2I is infeasible due to differing input modalities, yet our image-guided approach provides more detailed and robust attribute binding. Crucially, compared to existing image-guided techniques, we lead in both layout fidelity and LayoutSam-Eval color/texture accuracy.
Qualitative analysis, as presented in Fig. 5, highlights two key capabilities of our method. First, our approach effectively handles instance overlap, a common challenge for existing methods which often leads to attribute leakage or instance missing/merging. Second, our method exhibits superior spatial layout control, allowing it to synthesize a coherent and well-structured image from source images that may lack consistency. Additionally, as demonstrated in Fig. 6, our method performs robustly on complex text prompts, accurately reflecting fine-grained textual details in the generated image while preserving precise layout control.
4.4 Ablation Study
Table 3: Ablation study on applying ICA to different DiT-Blocks. $F$, $M$, $B$ denote FR-19, MID-19, BK-19 blocks respectively. Gray line denotes the method w/o CLA.
$F$ $M$ $B$ ITC AES IDS IPS AVG
83.16 53.80 22.70 72.45 58.03
91.54 58.41 24.19 74.46 62.15
91.14 57.36 26.08 74.17 62.19
91.42 57.76 26.57 74.39 62.53
91.20 57.00 30.80 77.63 64.16
90.26 58.35 31.27 76.99 64.22
91.55 58.85 31.10 75.64 64.28
91.38 58.24 32.72 76.32 64.66
Attention Mechanism Variations Across DiT-Blocks
We perform an ablation study to investigate the contribution of the ICA mechanism within our hierarchical attention architecture. We empirically divide the 57 DiT-blocks into three groups: FR-19 (first 19 blocks), MID-19 (middle 19 blocks), and BK-19 (last 19 blocks). The quantitative results on LAMIC-Bench++ are summarized in Tab. 3.
Prior work (Zhou et al., 2025) has demonstrated that MID-19 blocks have the most significant influence on instance-specific attributes. In alignment with this finding, our experiments confirm that applying the ICA mechanism selectively to the MID-19 blocks yields the highest average score of 64.66 and the best IDS score of 32.72. This configuration significantly outperforms the baseline that only uses the CLA mechanism, highlighting that targeted application of ICA is crucial for enhancing identity preservation and overall performance.
DPO Fine-tuning Analysis
To mitigate the model’s tendency to rigidly copy layout images while neglecting instance adaptation (e.g., posture, lighting), we employ Direct Preference Optimization (DPO) (Rafailov et al., 2024) with target images as preferred samples and layout images as less preferred. LoRA fine-tuning (Rank 256) is conducted with varying $\beta$ coefficients.
Table 4: Ablation study on DPO $\beta$.
DPO $\beta$ ITC AES IDS IPS AVG
100 91.32 57.97 22.36 74.54 61.55
250 91.44 57.58 24.49 75.01 62.13
500 91.33 57.57 25.01 74.92 62.21
750 91.13 57.22 25.89 75.45 62.42
1500 90.35 56.69 26.88 75.91 62.45
w/o DPO 86.84 54.19 32.37 76.78 62.55
1000 91.03 57.10 26.83 75.71 62.67
Tab. 4 and qualitative results in Sec. A.3 reveals three key findings:
  • Improved Composition: ITC and AES increase by $\uparrow$4.19–4.60% and $\uparrow$3.78–3.91% respectively across all $\beta$, demonstrating enhanced scene understanding.
  • Controlled Trade-off: IDS and IPS show moderate decreases ($\downarrow$6.49% and $\downarrow$1.07% at $\beta=1000$), with degradation scaling monotonically with $\beta$.
  • Optimal Configuration: $\beta=1000$ achieves the best balance (AVG 62.67), surpassing both non-DPO baseline (62.55) and other $\beta$ values.
This validates DPO’s ability to navigate the layout-adaptation trade-off.
5 Conclusion
In this work, we presented ContextGen, a novel framework for multi-instance generation that achieves precise control over layout and identity. Our approach is built on a unified token sequence that integrates text, layout, and multiple reference images, enabling a comprehensive understanding of the generation task. We introduced two core components: the Contextual Layout Anchoring (CLA) mechanism for enforcing robust spatial structure and the Identity Consistency Attention (ICA) mechanism for preserving fine-grained instance-specific attributes. Furthermore, our hierarchical attention architecture effectively leverages the intrinsic functional specialization of a Diffusion Transformer, with different layers dynamically attending to global and local contexts. To facilitate future research in this area, we created and will release the first large-scale, hierarchically-structured dataset, IMIG-100K, which we used to demonstrate our method’s superiority. Our extensive qualitative and quantitative evaluations show that ContextGen consistently outperforms state-of-the-art models, proving the efficacy of our design. We believe that this work provides a new foundation for the development of highly controllable and scalable multi-instance generation systems.
Appendix A More Implementation Details
A.1 Base Model Selection
In our framework, the selection of the base diffusion model is crucial for achieving high-fidelity multi-instance generation. We evaluated three variants of the FLUX family of models as potential backbones: FLUX.1-Dev (a general image generation model) (Labs, 2024b), FLUX.1-Fill (a local inpainting model) (Labs, 2024a), and FLUX.1-Kontext (an editing model) (Labs et al., 2025). While existing multi-subject-driven generation methods without layout control (Wu et al., 2025c; Hu et al., 2025) that do not rely on attention masks have successfully utilized FLUX.1-Dev, our experiments showed a significant limitation: without additional fine-tuning, FLUX.1-Dev failed to produce coherent images when an attention mask was applied. In contrast, both FLUX.1-Fill and FLUX.1-Kontext demonstrated the ability to generate images correctly with the attention mask. Among these two, FLUX.1-Kontext exhibited a noticeably superior capacity for identity preservation. Therefore, we chose FLUX.1-Kontext as the foundational model for our framework, leveraging its robust out-of-the-box performance with attention masking and its strong identity preservation capabilities.
A.2 Details of Compositing Layout Image
Our Contextual Layout Anchoring (CLA) mechanism relies on a meticulously constructed composite layout image to achieve robust spatial control. This process involves two key steps: determining the optimal composition order for all instances and then precisely placing each instance onto the canvas.
A correct composition order is crucial for multi-instance synthesis, especially when handling occlusions and complex overlaps. We propose a dynamic sorting algorithm, Instance Layering Prioritization, which first handles explicit containment relationships by prioritizing instances whose masks are completely contained within another’s. For all other candidate instances, we use an innovative hybrid priority scoring system to simulate the natural layering of objects. We utilize a pre-processing step to obtain each instance’s precise effective area (Ravi et al., 2024). The priority score $P_{i}$ is calculated as:
$P_{i}=\alpha\cdot\mathcal{A}(\text{instance}_{i})+\beta\cdot\left(1-\sum_{j\neq i}\text{IoU}(\text{instance}_{i},\text{instance}_{j})\right)+\lambda\cdot\text{RandomFactor}$
where $\mathcal{A}(\text{instance}_{i})$ is the area of instance $i$, $\text{IoU}(\text{instance}_{i},\text{instance}_{j})$ is the Intersection over Union between instances, and $\alpha,\beta,\lambda$ are hyperparameters.
The proposed hybrid priority scoring system is designed to simulate general, high-probability composition orders for model training. The introduction of the random factor enhances data diversity and model robustness during training. Despite the supplementary Identity Consistency Anchoring (ICA) mechanism, the overlap relationships can still influence the final generated image. Thus, for inference, a user-provided layout offers a more direct and customized form of control.
A.3 Details of DPO Fine-tuning
The visualization in Fig. 7 illustrates the efficacy of Direct Preference Optimization (DPO) in enhancing image generation, particularly by mitigating the issue of rigidly copying the layout image with blank backgrounds. To demonstrate this, we intentionally select an input and seed configuration that typically results in a minimal background scene.
The initial result on the left of Fig. 7(a) correctly anchors the main subject but suffers from the blank background issue, failing to render any environmental context. As the fine-tuning process advances, the model begins to introduce more naturalistic details, first by generating a realistic shadow and subsequently by adding a simple yet coherent background. Upon convergence, the final image on the right features a rich, detailed background, effectively demonstrating DPO’s ability to enrich the overall scene while strictly preserving the subject’s layout.
Fig. 7(b) shows how the DPO $\beta$ parameter affects the generation quality. A high $\beta$ value may limit the model’s capacity for meaningful tuning, whereas an overly low $\beta$ value can cause the model to follow the preference data too aggressively, risking a loss of the subject’s identity during convergence (as seen with the leather bag when $\beta=50$). Our results validate that when $\beta$ is correctly calibrated, DPO substantially improves image quality with minimal compromise to the subject’s identity.
(a) Results during the DPO fine-tuning process ($\beta=100$)
(b) Results at fine-tuning convergence with different DPO $\beta$
Figure 7: Results for the same input and fixed seed of DPO fine-tuning.
Appendix B More Experiment Details
B.1 Experimental Details on LAMICBench++
As shown in Tab. 1, we benchmark our method against several strong baselines, including single-image-editing models (Wu et al., 2025a; Labs et al., 2025) and closed-source commercial models (OpenAI, 2024; DeepMind, 2025). Since these two categories of models require different input modalities and instruction methods, we prepared distinct inputs for a fair evaluation:
Single-Image-Editing Models
These models are primarily designed for single-image editing, meaning they must process all instances combined within a single input image. We thus provided our manually composited layout images, ensuring minimal overlap across instances to avoid ambiguity. The following prompt template was used: ”Use the objects or humans in the image to create a new image that shows ’{PROMPT}’. Preserve object features and human identities (if any, including facial details). You may fill in the background with appropriate details to achieve a natural and aesthetically harmonious result.
Closed-Source Commercial Models
These are general-purpose multi-modal models that accept multiple input images. To guide them to perform the specific multi-instance generation task while preserving identities, we relied on a dedicated prompt alongside the references. The prompt template was: ”Generate a high-quality image of ’{PROMPT}’. Use the provided references, preserve object features and human identities (if any, including facial details).
B.2 More Qualitative Results on LAMICBench++
More qualitative results on the LAMICBench++ benchmark, as shown in Fig. 8, further validate our method’s superior performance, particularly in preserving subject identity. For example, our model successfully reconstructs the reference individuals in the first two examples with high fidelity in new contexts. Our framework’s strength is most apparent in its ability to handle complex compositional tasks. In the third example, the task requires generating multiple entities in different styles (Anime, realistic, etc.), and integrating them into a unified scene. Our model not only accurately preserves all subjects identities, but also cohesively integrates them into the stylized output. This highlights the robust scalability of our framework to handle intricate compositional tasks that combine multiple identities and styles.
B.3 More Results on COCO-MIG
Full quantitative result on COCO-MIG benchmark is shown in Tab. 5. Our method establishes a new state-of-the-art on all key metrics, achieving the highest average Success Rate (33.12%) and mIoU (65.12%) among all compared methods. Notably, our performance advantage is most pronounced in complex, high-instance-count scenarios. For $L_{4}$, $L_{5}$, and $L_{6}$ levels, our method significantly outperforms all baselines with a Success Rate of 28.12%, 23.12%, and 24.38%, respectively. This demonstrates the robust scalability of our hierarchical architecture to maintain both layout and identity control in intricate scenes. While some competitors show slightly higher scores on individual metrics like Global Clip (CreatiLayout (Wu et al., 2025c)) or Local Clip (InstanceDiffusion (Wang et al., 2024)), our approach achieves a superior overall balance. The consistently high mIoU scores across all complexity levels and our leading Instance Success Rate on the most challenging cases further validate our model’s ability to master both precise instance placement and high-fidelity attribute preservation.
More qualitative results on COCO-MIG are shown in Fig. 9. Our method demonstrates a clear qualitative advantage over existing models on the COCO-MIG benchmark. In the first example, other methods fail to correctly generate the blue vase, with issues ranging from incorrect position to instance merging. Our model, in contrast, precisely renders the vase as intended. Similarly, for the ’green potted plant’, our approach correctly applies the ’green’ attribute to the pot, whereas competitors fail to do so. This highlights our superior ability to handle holistic subject identity. Furthermore, while some baselines correctly generate the requested objects in the third and fourth examples, their outputs often lack aesthetic harmony and visual coherence. Our method consistently produces images that are not only accurate but also visually pleasing and well-composed. These results underscore two key advantages of our framework: (1) Compared to other image-guided methods like MS-Diffusion (Wang et al., 2025), our approach offers significant superiority in layout control (2) Our dedicated identity preservation mechanism provides more robust and reliable subject fidelity than the attribute-based control of text-guided methods, particularly in intricate, multi-instance scenes.
B.4 More Qualitative Results on LayoutSam-Eval
Additional qualitative results are presented in Fig. 10. Evidently, our method exhibits superior overall visual quality and realism compared to all existing approaches. While other methods (especially those reliant on text-guided layout-to-image generation) struggle with preserving fine-grained attributes—such as the specific text on the building in the first example and the exact color of the man’s shorts in the second—our approach faithfully preserves the user’s intended details to the greatest extent, excelling in both layout control and attribute binding.
Table 5: Full quantitative result on COCO-MIG. According to the count of generated instances, COCO-MIG is divided into five levels: $L_{2}$, $L_{3}$, $L_{4}$, $L_{5}$, and $L_{6}$. $L_{i}$ means that the count of instances needed to generate in the image is ii.
Figure 8: More qualitative results on LAMICBench++.
Figure 9: More qualitative results on COCO-MIG.
Figure 10: More qualitative results on LayoutSam-Eval.
Appendix C Limitations and Future Work
While our framework demonstrates state-of-the-art performance in multi-instance generation, it is not without limitations. A primary challenge stems from our model’s strong emphasis on identity preservation. When inconsistencies exist between the provided reference images or between the images and the text prompt, our model tends to prioritize maintaining the identities of the reference subjects. This can sometimes lead to a lack of flexibility in adjusting attributes such as lighting, color, or pose, which may compromise the overall visual harmony and text-image consistency of the final output. This trade-off between identity fidelity and contextual flexibility represents an important area for future research. In the future, we plan to explore more dynamic attention mechanisms that can better balance these competing demands, allowing for more flexible style and attribute transfer while preserving core subject identities.
Appendix D The Use Of Large Language Models
During the preparation of this manuscript, Large Language Models were used as a general-purpose writing assistant tool. Specifically, LLMs were employed to polish the language and refine the clarity of the text. The authors take full responsibility for the content of the paper.